how do you find the zeros in this equation?
x^5-5x^4-3x^3+15x^2-4x+20=0

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Try the factors of 20, plus and minus
Sub each for x and see if it's zero.

Answer by stanbon(75887)

find the zeros in this equation?
x^5-5x^4-3x^3+15x^2-4x+20=0
---
I graphed it and found Real zeros at x = 2 and x = -2 and x = 5
----
Use synthetic division to look for other zeros:
2)....1....-5....-3....15....-4....20
........1.....-3....-9....-3....-10...|..0
====
-2)......1......-5....1.....-5....|..0
====
5)........1.......0.....1.....|0
----
Quotient: x^2+1
---
Complex zeros:
x^2+1 = 0
(x+i)(x-i) = 0
x = -i or x = i
=========================
Cheers,
Stan H.